OpenStudy (anonymous):

For y = x2 + 6x − 16, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation. Show all work and use complete sentences

OpenStudy (jhannybean):

We can find the vertex first \[y=x^2+6x-16\]\[y+16=x^2+6x\] now we'll be completing the square . \[y+16=x^2+\color{red}6x\]\[y+16+\color{red}9=x^2+6x+\color{red}{9}\]\[y+25=(x+3)^2\]\[y=(x+3)^2-25\]\[y=(x-(\color{blue}{-3}))+(\color{blue}{-25})\] Your vertex.

OpenStudy (luigi0210):

we know from the previous it ups upward

OpenStudy (luigi0210):

*opens

OpenStudy (luigi0210):

x intercept when y equals 0

OpenStudy (primeralph):

it's simple. If the co-eff of x^2 is positive, it opens up, if negative, opens down.

OpenStudy (luigi0210):

we could use the first derivative test to determine if it's a max or min.. since it opens up it's a min..

OpenStudy (jhannybean):

The formula \[\large y=\color{purple}a(x-h)^2+k\] is the formula for a quadratic, and your "a" value tells if it opens up or down, your formula \[\large y=\color{purple}{(1)}(x-(-3))^2-25\] the "a" is positive, so your parabola opens upward.

OpenStudy (primeralph):

@Jhannybean good job.

OpenStudy (luigi0210):

It's her thing :P

OpenStudy (jhannybean):

you can find your x-intercept by setting y=0. \[\large (x+3)^2-25=0\]\[\large (x+3)^2=25\]\[\large \sqrt{(x+3)^2}= \pm \sqrt{25}\]\[\large |x+3|= \pm \sqrt{25}\] solving for x would give you your x-intercepts. \[\large |x+3| = \pm 5\]

OpenStudy (luigi0210):

0=(x+8)(x-2)

OpenStudy (jhannybean):

When describing the graph of the equation, you could use your min to state where the slope is inc/dec, and the domain/range of the function

OpenStudy (luigi0210):

@sumsumjoe we're not doing all this work for you if you're not gonna try and understand..

OpenStudy (luigi0210):

Well Jhan won't :P

OpenStudy (jhannybean):

^_^

OpenStudy (luigi0210):

and as always.. great job :D

OpenStudy (jhannybean):

Lol, thanks

OpenStudy (luigi0210):

I think he's gone :l